OpenStudy (dls):

Password of lock is a 5-digit number. A little information about it is as follows: · The fourth digit is 4 greater than the second digit. · The third digit is 3 less than the second digit. · The first digit is thrice the last digit. · There are 3 pairs of digits whose sum is 11. Password=?

OpenStudy (dls):

Can't we let the 5 digits be abcde and then..solve..like d=4+b c=3-b.. c+d=7...etc?

OpenStudy (anonymous):

abcde according the question, d=b+4 c=b-3 a=3e So, the lock becomes (3e)b(b-3)(b+4)e

OpenStudy (dls):

haha!

OpenStudy (anonymous):

Also, 3<=b<=5

OpenStudy (anonymous):

0<=e<=3

OpenStudy (dls):

d=4+b c=b-3 this gives..d-c=7

OpenStudy (dls):

How to utilize this? There are 3 pairs of digits whose sum is 11.

OpenStudy (shubhamsrg):

since 2nd digit can only have values from 3 and 5 for c and d to have values between 0 and 9 , we see there are only 3 possible combos for 2nd, 3rd and 4th digits. (3, 0, 7) (4, 1, 8) (5, 2, 9) It is given that there are 3 pairs whose sum is 11. 11= 2+ 9= 3+ 8=4+ 7=5+ 6. so we can not have a 0 or 1 there.. hence the ans must be _529_ 1st digit=3* last digit. The possible combinations are (3, 1), (6, 2) and (9, 3) hit and trial shows only 6 and 2 fir it for 3 pairs summing to 11 hence ans=65292. please confirm..

OpenStudy (dls):

yes its correct

OpenStudy (shubhamsrg):

phew!

OpenStudy (dls):

bingo!